OpenStudy (djbreezy):

Complete the following steps for this equation. 5x – 20 = 4y 1. Write the equation in standard form. 2. Find the x-intercept. 3. Find the y-intercept. 4. Graph the linear equation on the coordinate plane.

563blackghost (563blackghost):

To change to standard form you would add \(\Large{20}\) to the left side to cancel it out and add it to the other.... \(\LARGE{5x(−20+20)=4y+20→5x=4y+20}\) In which you would subtract \(\Large{4y}\) from the right side and add the \(\Large{−4y}\) to the other side....so the standard equation would look like this.... \(\Huge\color{purple}{5x−4y=20}\) Do you understand my process?

563blackghost (563blackghost):

are u there @DJBreezy

OpenStudy (djbreezy):

Sorry.....im back

OpenStudy (djbreezy):

@563blackghost ..can u show me how to do it give all the work and the answer so therefore next time i know how to do it next time i end up doing a question like this?

563blackghost (563blackghost):

Well I can't do that since that is basically given away the answer....

563blackghost (563blackghost):

To find the x-intercept you would basically make y as 0 so you would input that and solve for the x-intercept.... \(\Huge{5x-4(0)=20}\) Simplify to solve... To find the y-intercept you would follow the same process EXCEPT you would make x as 0 instead of y.... \(\Huge{5(0)-4y=20}\)